A school survey found that 9 out of 10 students like pizza. If three students are chosen at random with replacement, what is the probability that all the three students like pizza?


1
Expert's answer
2021-06-15T13:45:27-0400

Binomial distribution

"P=\frac{9}{10}=0.9"

"n=3"

"P(x)=\binom{n}{x}p^x(1-p)^{n-x}, x=0,1,2,3"

"P(3)=\binom{3}{3}0.9^3=0.729"